# Look, Listen, and Answer: Overcoming Biases for Audio-Visual Question Answering

### v1

@@ 5 Experiments
Type Method Audio QA Visual QA AVQA All
@@
Audio QA FCNLSTM 70.45 66.22 68.88 63.89 46.74 55.21 82.01 46.28 59.34 62.15 47.33 60.06 60.34
CONVLSTM 74.07 68.89 72.15 67.47 54.56 60.94 82.91 50.81 63.03 60.27 51.58 62.24 63.65
Visual QA BiLSTM Attn 70.35 47.92 62.05 64.64 64.33 64.48 78.39 45.85 56.91 53.09 49.76 57.10 59.92
HCAttn 70.25 54.91 64.57 64.05 66.37 65.22 79.10 49.51 59.97 55.25 56.43 60.19 62.30
MCAN 77.50 55.24 69.25 71.56 70.93 71.24 80.40 54.48 64.91 57.22 47.57 61.58 65.49
GRU 72.21 66.89 70.24 67.72 70.11 68.93 81.71 59.44 62.64 61.88 60.07 65.18 67.07
Video QA HCRN 68.59 50.92 62.05 64.39 61.81 63.08 54.47 41.53 53.38 52.11 47.69 50.26 55.73
HME 74.76 63.56 70.61 67.97 69.46 68.76 80.30 53.18 63.19 62.69 59.83 64.05 66.45
PSAC 75.64 66.06 72.09 68.64 69.79 69.22 77.59 55.02 63.42 61.17 59.47 63.52 66.54
AVQA AVSD 72.41 61.90 68.52 67.39 74.19 70.83 81.61 58.79 63.89 61.52 61.41 65.49 67.44
LAViT 74.36 64.56 70.73 69.39 75.65 72.56 81.21 59.33 64.91 64.22 63.23 66.64 68.93
STG 78.18 67.05 74.06 71.56 76.38 74.00 81.81 64.51 70.80 66.01 63.23 69.54 71.52
COCA 79.35 66.50 74.61 72.35 76.08 74.24 83.50 64.02 70.99 63.40 64.48 69.47 71.64
Ours 83.87 71.04 79.14 79.78 76.73 78.24 80.87 51.63 71.46 64.67 64.60 67.13 72.20
5.1 Dataset and Evaluation
MUSIC-AVQA Li et al. (2022a), which contains training, validation, and testing splits with 31,927, 4,568, and 9,129 QA pairs, is developed by gathering questions for 9,288 musical performances. The questions are produced by a limited set of pre-defined templates. The videos, sourced from YouTube, include solo performances, ensembles of the same instruments, and ensembles of different instruments. This dataset consists of three tasks: audio QA, visual QA, and AVQA. The standard accuracy is used to evaluate model performance on the mentioned tasks. To comprehensively assess model robustness, we conduct rephrasing and splitting on the test split, expanding the question count from 9,129 to 211,572. The necessary discussion on data annotations is shown in Appendix D.1. Owing to the introduction of distribution shift, our proposed dataset provides three metrics: head accuracy, tail accuracy, and overall accuracy, to evaluate models precisely. The test split comparison between MUSIC-AVQA and MUSIC-AVQA-R is shown in Appendix D.2. We can see that the latter exhibits a larger test sample space.
Method Audio QA Visual QA AVQA All
CNT COMP CNT LOC EXIST LOC CNT COMP TEMP Avg.
H T H T H T H T H T H T H T H T H T
FCNLSTM 66.23 36.48 64.78 51.14 61.75 5.31 54.86 51.06 64.76 78.52 46.66 57.30 62.69 7.23 43.13 71.67 37.02 30.78 54.12
CONVLSTM 70.22 41.14 67.50 52.93 62.11 9.17 53.44 49.88 60.08 84.82 46.46 59.90 56.52 8.18 43.29 72.52 41.54 45.12 55.20
BiLSTM Attn 73.68 46.32 21.51 77.58 64.30 0.00 53.92 42.01 87.51 21.14 35.16 43.75 62.85 2.18 27.61 74.38 17.58 31.32 48.84
HCAttn 61.67 41.63 59.09 47.14 56.52 9.20 67.01 53.16 66.57 61.13 37.05 42.48 59.53 12.48 48.81 60.12 33.82 39.26 51.90
GRU 66.92 48.63 58.29 59.61 64.37 11.79 57.68 57.66 76.30 64.76 41.05 45.61 60.71 18.68 57.19 57.38 31.02 40.67 55.21
MCAN 75.02 60.16 58.89 50.09 64.58 26.69 66.48 62.25 51.29 67.29 46.11 61.61 64.76 25.28 50.57 52.40 34.64 58.05 57.27
HCRN 55.53 53.31 47.17 32.44 41.87 23.55 39.40 51.27 41.81 65.45 36.62 42.72 54.58 19.57 33.33 36.87 40.47 44.13 43.92
HME 62.60 53.95 54.97 58.29 50.95 16.46 73.25 58.60 65.74 66.49 33.79 46.03 63.18 17.18 53.20 60.57 33.95 41.57 53.66
PSAC 53.01 56.68 57.41 48.12 49.55 26.43 72.96 60.69 50.56 55.54 41.98 52.30 56.70 19.58 38.13 58.92 26.68 46.24 50.45
AVSD 54.00 47.84 60.61 47.79 60.34 10.07 74.78 61.43 66.28 61.98 33.00 40.35 46.21 8.06 51.98 66.00 40.14 41.52 52.33
LAViT 50.57 43.45 50.78 44.93 47.28 15.50 67.19 65.51 52.37 22.04 44.35 61.69 52.21 21.52 45.61 40.49 35.00 49.33 47.40
STG 56.40 41.48 62.28 57.59 59.86 12.94 64.31 54.00 73.35 77.26 35.35 40.49 48.31 8.41 53.30 62.44 40.25 38.15 52.80
Ours 84.32 67.23 64.68 62.18 75.09 48.42 80.47 66.38 77.22 67.58 55.15 82.23 70.12 39.83 61.26 58.17 43.67 58.33 66.95
@@ 5.3 Baselines
We select 13 previous state-of-the-art multi-modal QA methods as baselines to verify the effectiveness of the proposed architecture and investigate the robustness of these methods. Audio QA methods: FCNLSTM Fayek and Johnson (2020), and CONVLSTM Fayek and Johnson (2020). Visual QA methods: GRU Antol et al. (2015), BiLSTM Attn Zhou et al. (2016), HCAttn Lu et al. (2016), and MCAN Yu et al. (2019). Video QA methods: HME Fan et al. (2019), PSAC Li et al. (2019b), and HCRN Le et al. (2020). AVQA methods: AVSD Schwartz et al. (2019), LAViT Yun et al. (2021), STG Li et al. (2022a), and COCA Lao et al. (2023). For fair comparisons, we do not select large models as backbones for our architecture. Additionally, we abstain from reassessing COCA on the MUSIC-AVQA-R dataset due to its lack of publicly available code. The baseline introductions are shown in Appendix D.4.
@@ 5.4 Comparison on MUSIC-AVQA
We conduct experiments on the MUSIC-AVQA test split to validate the effectiveness of the proposed architecture. The results are presented in Table 1. Notably, our architecture obtains significant improvements of 4.53% and 4% in the audio and visual QA tasks, respectively. It also achieves the best performance in the AVQA task. Furthermore, our architecture obtains a new state-of-the-art result of 72.20% on the whole question. It is worth mentioning that visual QA methods, like GRU, and video QA methods, such as HME, also exhibit competitive results in the AVQA task, despite lacking one modality as input. The results on this dataset can demonstrate the efficacy of these methods to some extent. However, MUSIC-AVQA lacks refined and precise evaluations due to its inherent shortcomings that are introduced in Section 1. Consequently, it is insufficient to evaluate these methods only on this dataset.
@@ 5.7 Sensitivity and Qualitative Analysis
6 Conclusion
We are the first to investigate bias learning in the AVQA task from model evaluation and design aspects. On the one hand, we construct a new dataset, MUSIC-AVQA-R, which evaluates the performance on the head, tail, and overall samples, providing a precise measure of model robustness. On the other hand, we introduce a robust architecture employing the MCCD strategy to mitigate bias learning. Extensive experiments demonstrate the effectiveness of our architecture and the debiasing capability of MCCD. Furthermore, we reevaluate previous multi-modal QA methods on our proposed dataset, revealing their poor robustness.
Limitations
We explore potential limitations not only in terms of model measurement but also in model designs. Although MUSIC-AVQA-R is the first AVQA dataset focusing on robustness evaluations, it still has shortcomings in the answer annotation. Due to constraints imposed by MUSIC-AVQA, the answer space is limited, comprising only 42 classes, and answer lengths are typically confined to a single word. This deviation from real-world scenarios is noteworthy. Concerning model designs, for a fair comparison with baselines, like LAViT using a pre-trained model, we do not select large generative models to be backbones. However, compared with the answer classification, it may be more useful to generate answers for the AVQA task. Furthermore, our model only provides attention weights of input modalities for humans, lacking the capability of generating sonically and visually grounded responses. This capability is potentially crucial for facilitating humans’ intuitive comprehension of the answers.
Ethics Statement
MUSIC-AVQA Li et al. (2022a) has undergone meticulous pre-processing tailored for academic research purposes. We ensure the absence of any information that discloses the names or uniquely identifies individuals, as well as avoiding any offensive content. We only perform rephrasing and splitting for the questions within the dataset to develop MUSIC-AVQA-R, which preserves its inherent characteristics.
@@ Appendix B Dataset Analysis
We visualize the answer distribution of specific types of questions. Fig. 6, 7, and 8 show the visualization of the AVQA, audio QA and visual QA tasks, respectively. We can see that all of the answer distributions are long-tail. This further demonstrates the necessity of head and tail sample splitting. Specifically, for the “Location” type in the AVQA task, we can see that the number of “congas” is far less than that of “yes”. It is noteworthy that, for question types featuring only two possible answers, we classify questions with lower frequencies as tail samples and those with higher frequencies as head samples. For instance, for the “Existential” questions in the AVQA task, we consider the questions with the answer “no” as tail samples.

### v3

@@ 5 Experiments
5.1 Dataset and Evaluation
MUSIC-AVQA [4], which contains training, validation, and testing splits with 31,927, 4,568, and 9,129 QA pairs, is developed by gathering questions for 9,288 musical performances. The questions are produced by a limited set of pre-defined templates. The videos, sourced from YouTube, include solo performances, ensembles of the same instruments, and ensembles of different instruments. This dataset consists of three tasks: audio QA, visual QA, and AVQA. The standard accuracy is used to evaluate model performance on the mentioned tasks. To comprehensively assess model robustness, we conduct rephrasing and splitting on the test split, expanding the question count from 9,129 to 211,572. Owing to the introduction of distribution shift, our proposed dataset provides three metrics: head accuracy, tail accuracy, and overall accuracy, to evaluate models precisely. The test split comparison between MUSIC-AVQA and MUSIC-AVQA-R is shown in Appendix D.1. We can see that the latter exhibits a larger test sample space.
Method MCCD Audio QA Visual QA AVQA All
@@ 5.1 Dataset and Evaluation
FCNLSTM $\times$ 70.45 66.22 68.88 63.89 46.74 55.21 82.01 46.28 59.34 62.15 47.33 60.06 60.34
$\checkmark$ 70.99 66.50 69.34 66.08 59.02 62.51 83.50 57.17 60.47 61.58 57.54 64.11 64.61
CONVLSTM $\times$ 74.07 68.89 72.15 67.47 54.56 60.94 82.91 50.81 63.03 60.27 51.58 62.24 63.65
$\checkmark$ 72.76 69.53 71.57 69.59 58.12 63.79 82.69 56.09 62.13 62.03 55.11 63.87 65.21
BiLSTM Attn $\times$ 70.35 47.92 62.05 64.64 64.33 64.48 78.39 45.85 56.91 53.09 49.76 57.10 59.92
$\checkmark$ 68.24 54.88 63.31 61.65 55.92 58.75 79.15 41.96 55.02 49.41 49.15 55.18 57.56
HCAttn $\times$ 70.25 54.91 64.57 64.05 66.37 65.22 79.10 49.51 59.97 55.25 56.43 60.19 62.30
$\checkmark$ 69.52 53.37 63.56 63.99 65.47 64.74 78.64 47.28 61.11 55.86 55.72 60.03 61.90
MCAN $\times$ 77.50 55.24 69.25 71.56 70.93 71.24 80.40 54.48 64.91 57.22 47.57 61.58 65.49
$\checkmark$ 78.27 56.57 70.27 71.93 71.18 71.55 81.48 54.24 65.77 55.86 46.84 61.54 65.74
GRU $\times$ 72.21 66.89 70.24 67.72 70.11 68.93 81.71 59.44 62.64 61.88 60.07 65.18 67.07
$\checkmark$ 73.35 66.16 70.70 67.25 71.43 69.36 81.98 60.11 63.08 62.76 61.19 65.84 67.63
HCRN $\times$ 68.59 50.92 62.05 64.39 61.81 63.08 54.47 41.53 53.38 52.11 47.69 50.26 55.73
$\checkmark$ 72.17 64.65 69.40 67.42 60.82 64.08 79.66 48.70 65.14 61.22 55.72 60.40 64.20
HME $\times$ 74.76 63.56 70.61 67.97 69.46 68.76 80.30 53.18 63.19 62.69 59.83 64.05 66.45
$\checkmark$ 72.96 62.29 69.03 68.76 69.31 69.03 80.77 52.61 62.92 63.03 60.71 64.19 66.33
PSAC $\times$ 75.64 66.06 72.09 68.64 69.79 69.22 77.59 55.02 63.42 61.17 59.47 63.52 66.54
$\checkmark$ 75.02 65.66 71.57 69.09 69.88 69.49 79.35 53.04 61.98 61.13 57.66 62.85 66.15
AVSD $\times$ 72.41 61.90 68.52 67.39 74.19 70.83 81.61 58.79 63.89 61.52 61.41 65.49 67.44
$\checkmark$ 72.07 63.97 69.09 67.42 74.53 71.02 81.17 59.13 63.08 62.49 63.50 65.82 67.77
LAViT $\times$ 74.36 64.56 70.73 69.39 75.65 72.56 81.21 59.33 64.91 64.22 63.23 66.64 68.93
$\checkmark$ 75.12 65.49 71.57 70.43 76.73 73.62 81.38 60.33 65.30 62.49 62.29 66.42 69.24
STG $\times$ 78.18 67.05 74.06 71.56 76.38 74.00 81.81 64.51 70.80 66.01 63.23 69.54 71.52
COCA $\times$ 79.35 66.50 74.61 72.35 76.08 74.24 83.50 64.02 70.99 63.40 64.48 69.47 71.64
Ours $\checkmark$ 83.87 71.04 79.14 79.78 76.73 78.24 80.87 51.63 71.46 64.67 64.60 67.13 72.20
LAVisH $\times$ 81.32 63.30 74.67 79.20 80.57 79.89 83.40 65.22 72.96 64.03 66.18 70.57 73.76
$\checkmark$ 80.33 63.80 74.24 78.86 81.31 80.10 73.91 65.22 73.91 64.31 66.55 70.80 73.87
@@ 5.3 Baselines
We select 14 previous state-of-the-art multi-modal QA methods as baselines to verify the effectiveness of the proposed architecture and investigate the robustness of these methods. Audio QA methods: FCNLSTM [56], and CONVLSTM [56]. Visual QA methods: GRU [57], BiLSTM Attn [58], HCAttn [59], and MCAN [60]. Video QA methods: HME [61], PSAC [62], and HCRN [63]. AVQA methods: AVSD [13], LAViT [5], STG [4], COCA [53] and LAVisH [1]. We abstain from reassessing COCA on the MUSIC-AVQA-R dataset due to its lack of publicly available code. The baseline introductions are shown in Appendix D.3. Due to the particularly slow computation speed of STG, we do not conduct experiments with STG+MCCD. Due to computing power limitations, we reevaluate LAVisH with a batch size of 2.
@@ 5.4 Comparison on MUSIC-AVQA
We conduct experiments on the MUSIC-AVQA test split to validate the effectiveness of the proposed architecture. The results are presented in Table 1. All methods, except LAVisH, utilize ResNet-18 to encode visual features, whereas LAVisH employs stronger models such as ViT [64] or Swin [65] to acquire visual representations. We first analyze the comparison under the same visual encoding conditions. Notably, compared with COCA, our architecture obtains significant improvements of 4.53% and 4% in the audio and visual QA tasks, respectively. It also achieves the best performance in the AVQA task. Furthermore, our architecture obtains a new state-of-the-art result of 72.20% on the whole question. It is worth mentioning that visual QA methods, like GRU, and video QA methods, such as HME, also exhibit competitive results in the AVQA task, despite lacking one modality as input. We also observe that LAVisH, proposed based on STG, introduces trainable parameters into robust visual encoders, thereby achieving superior results compared to methods employing weaker visual encoders. The results on this dataset can demonstrate the efficacy of these methods to some extent. However, MUSIC-AVQA lacks refined and precise evaluations due to its inherent shortcomings that are introduced in Section 1. Consequently, it is insufficient to evaluate these methods only on this dataset.
@@ 5.5 Robustness Evaluation
Method MCCD Audio QA Visual QA AVQA All
CNT COMP CNT LOC EXIST LOC CNT COMP TEMP Avg.
H T H T H T H T H T H T H T H T H T
FCNLSTM $\times$ 66.23 36.48 64.78 51.14 61.75 5.31 54.86 51.06 64.76 78.52 46.66 57.30 62.69 7.23 43.13 71.67 37.02 30.78 54.12
$\checkmark$ 62.51 34.44 61.19 51.26 61.11 5.66 57.73 50.36 62.48 82.40 45.49 60.09 62.07 7.16 44.55 69.46 36.55 30.74 54.55
CONVLSTM $\times$ 70.22 41.14 67.50 52.93 62.11 9.17 53.44 49.88 60.08 84.82 46.46 59.90 56.52 8.18 43.29 72.52 41.54 45.12 55.20
$\checkmark$ 68.38 41.58 68.39 52.10 61.46 9.56 54.17 50.33 59.61 83.11 55.29 56.52 59.13 7.82 45.31 72.70 41.26 45.40 55.74
BiLSTM Attn $\times$ 73.68 46.32 21.51 77.58 64.30 0.00 53.92 42.01 87.51 21.14 35.16 43.75 62.85 2.18 27.61 74.38 17.58 31.32 48.84
$\checkmark$ 73.30 45.16 20.71 77.48 64.41 0.00 56.08 42.54 87.47 21.04 34.47 43.51 63.33 2.18 26.01 75.48 17.92 32.67 49.55
HCAttn $\times$ 61.67 41.63 59.09 47.14 56.52 9.20 67.01 53.16 66.57 61.13 37.05 42.48 59.53 12.48 48.81 60.12 33.82 39.26 51.90
$\checkmark$ 62.50 41.43 58.89 47.42 56.65 8.85 67.31 52.92 66.82 59.87 38.25 42.53 59.38 12.42 57.39 52.01 32.84 39.55 52.29
GRU $\times$ 66.92 48.63 58.29 59.61 64.37 11.79 57.68 57.66 76.30 64.76 41.05 45.61 60.71 18.68 57.19 57.38 31.02 40.67 55.21
$\checkmark$ 69.94 48.09 56.31 63.77 66.24 13.36 63.55 57.59 83.04 54.16 43.36 43.36 57.89 18.36 53.93 59.65 30.82 38.23 55.70
MCAN $\times$ 75.02 60.16 58.89 50.09 64.58 26.69 66.48 62.25 51.29 67.29 46.11 61.61 64.76 25.28 50.57 52.40 34.64 58.05 57.27
$\checkmark$ 73.53 56.14 68.31 39.44 65.51 29.40 68.41 60.09 58.80 61.90 46.75 60.61 60.54 31.89 69.09 44.94 32.44 57.78 58.22
HCRN $\times$ 55.53 53.31 47.17 32.44 41.87 23.55 39.40 51.27 41.81 65.45 36.62 42.72 54.58 19.57 33.33 36.87 40.47 44.13 43.92
$\checkmark$ 51.96 49.21 43.42 36.78 41.13 20.71 37.79 50.99 44.38 58.40 35.05 46.33 54.39 20.90 34.50 33.14 40.13 44.00 42.87
HME $\times$ 62.60 53.95 54.97 58.29 50.95 16.46 73.25 58.60 65.74 66.49 33.79 46.03 63.18 17.18 53.20 60.57 33.95 41.57 53.66
$\checkmark$ 60.62 53.85 62.22 53.01 52.90 14.96 72.56 58.56 55.47 69.21 32.27 42.97 69.90 12.36 43.51 72.51 36.65 32.61 53.34
PSAC $\times$ 53.01 56.68 57.41 48.12 49.55 26.43 72.96 60.69 50.56 55.54 41.98 52.30 56.70 19.58 38.13 58.92 26.68 46.24 50.45
$\checkmark$ 55.14 52.26 64.70 44.45 52.34 22.15 72.06 60.70 58.97 52.35 41.18 49.78 53.28 18.85 42.60 64.53 25.81 45.68 51.31
AVSD $\times$ 54.00 47.84 60.61 47.79 60.34 10.07 74.78 61.43 66.28 61.98 33.00 40.35 46.21 8.06 51.98 66.00 40.14 41.52 52.33
$\checkmark$ 55.87 40.18 65.41 48.05 63.32 7.41 73.78 58.20 74.74 70.80 37.85 34.55 35.53 6.11 49.96 67.88 44.03 43.89 53.09
LAViT $\times$ 50.57 43.45 50.78 44.93 47.28 15.50 67.19 65.51 52.37 22.04 44.35 61.69 52.21 21.52 45.61 40.49 35.00 49.33 47.40
$\checkmark$ 45.05 45.09 57.33 41.26 48.62 17.00 69.91 65.90 60.61 29.57 43.17 57.57 53.92 22.09 54.46 35.35 33.99 49.40 48.91
STG $\times$ 56.40 41.48 62.28 57.59 59.86 12.94 64.31 54.00 73.35 77.26 35.35 40.49 48.31 8.41 53.30 62.44 40.25 38.15 52.80
LAVisH $\times$ 61.73 43.99 65.06 60.38 65.53 11.13 70.21 64.73 77.83 79.46 41.76 41.20 49.88 14.87 59.26 65.10 41.84 46.26 57.63
$\checkmark$ 74.02 65.17 64.73 53.15 71.96 40.56 68.49 66.00 63.17 66.68 30.11 43.80 63.77 26.51 56.31 63.46 50.79 42.85 59.25
Ours $\checkmark$ 84.32 67.23 64.68 62.18 75.09 48.42 80.47 66.38 77.22 67.58 55.15 82.23 70.12 39.83 61.26 58.17 43.67 58.33 66.95
To validate the plug-and-play capability of MCCD, we conduct extensive experiments using the baselines+MCCD on the aforementioned datasets. The results are presented in Tables 1 and 2. We observe that MCCD consistently improves performance across most methods on MUSIC-AVQA (9 out of 13) and MUSIC-AVQA-R (11 out of 13), respectively. This underscores the robust debiasing capability of MCCD in a plug-and-play manner.
@@ 5.6 Ablation Study
5.7 Sensitivity and Qualitative Analysis
6 Conclusion and Limitation
We are the first to investigate bias learning in the AVQA task from model evaluation and design aspects. On the one hand, we construct a new dataset, MUSIC-AVQA-R, which evaluates the performance on the head, tail, and overall samples, providing a precise measure of model robustness. On the other hand, we introduce a robust architecture employing the MCCD strategy to mitigate bias learning. Extensive experiments demonstrate the effectiveness of our architecture and the plug-and-play debiasing capability of MCCD. Furthermore, we reevaluate previous multi-modal QA methods on our proposed dataset, revealing their poor robustness.
Due to constraints imposed by MUSIC-AVQA, the answer space of our dataset is limited, comprising only 42 classes, and answer lengths are typically confined to a single word. This deviation from real-world scenarios is noteworthy. Concerning model designs, for a fair comparison with baselines, we do not select large generative models to be backbones. However, compared with the answer classification, it may be more useful to generate answers for the AVQA task.
7 Acknowledgements
This work was supported by the National Key Research and Development Program of China (2021YFB1715600), the National Natural Science Foundation of China (U22B2019, 62477037, 62450005, 62437002, 62306229, 62293553), the Natural Science Basic Research Program of Shaanxi (2023-JC-YB-593), the Youth Innovation Team of Shaanxi Universities “Multi-modal Data Mining and Fusion”, the Shaanxi Undergraduate and Higher Education Teaching Reform Research Program (23BY195), the Youth Talent Support Program of Shaanxi Science and Technology Association (20240113), the Xi’an Jiaotong University-China Mobile Communications Group Co., Ltd. Digital Government Joint Institute, and the China Postdoctoral Science Foundation (2024M752585).
@@ Appendix B Dataset Analysis
We make statistics on the rephrasings, as depicted in Table 4. Any rephrasing that garners fewer than one vote will be disregarded. It is evident that the overwhelming majority of the rephrased questions received three favorable votes.
Table 4: Statistics of rephrasing consistency. Positive and Negative denote whether the annotator agrees with the rephrasing or not.
Positive Negative Total
3 0 164, 219
2 1 47,353
1 2 7,481
0 3 9,172
Fig. 6 shows the distribution of questions in MUSIC-AVQA and MUSIC-AVQA-R based on the first three words. It is evident that there are notably more entries within each circle in the left figure compared to the right figure. This suggests that the diversity of our dataset is higher than MUSIC-AVA.
(a) MUSIC-AVQA-R.
(b) MUSIC-AVQA.
Figure 6: Distribution visualization of questions based on the first three words.
@@
We visualize the answer distribution of specific types of questions. Fig. 7, 8, and 9 show the visualization of the AVQA, audio QA and visual QA tasks, respectively. We can see that all of the answer distributions are long-tail. This further demonstrates the necessity of head and tail sample splitting. Specifically, for the “Location” type in the AVQA task, we can see that the number of “congas” is far less than that of “yes”. It is noteworthy that, for question types featuring only two possible answers, we classify questions with lower frequencies as tail samples and those with higher frequencies as head samples. For instance, for the “Existential” questions in the AVQA task, we consider the questions with the answer “no” as tail samples.
B.1 Ethics Statement
MUSIC-AVQA [4] has undergone meticulous pre-processing tailored for academic research purposes. We ensure the absence of any information that discloses the names or uniquely identifies individuals, as well as avoiding any offensive content. We only perform rephrasing and splitting for the questions within the dataset to develop MUSIC-AVQA-R, which preserves its inherent characteristics.
B.2 Question Comparison
We select questions from both the head and tail splits to demonstrate the diversity of our dataset, as illustrated in Figures 10 and 11. Due to the use of pre-defined templates, the questions in the train and test splits of MUSIC-AVQA differ by only a single word. In contrast, the questions in our dataset exhibit a variety of formats, which better reflect real-world scenarios. Additionally, our dataset encompasses a larger test space compared to MUSIC-AVQA.
Figure 10: Question Comparison between MUSIC-AVQA and MUSIC-AVQA-R. The rephrased question comes from the head split of our dataset. The questions in our dataset feature diverse formats, which more accurately reflect real-world scenarios.
Figure 11: Question Comparison between MUSIC-AVQA and MUSIC-AVQA-R. The rephrased question comes from the tail split of our dataset. The questions in our dataset feature diverse formats, which more accurately reflect real-world scenarios.